OpenStudy (anonymous):

if f(x)=x^2-3 and f(a)= 46, what is the value of a? Explain.

OpenStudy (aum):

\[ f(x)=x^2-3 \\ f(a) = a^2 - 3 = 46 \\ a^2 = 49 \\ a = \pm7 \]